solve: x-3/2 - 3x+4/2 =3
Answer by kwobsi(2)

x-3/2 - 3x+4/2 =3
multiply through the equation by 2
that is 2x-3-6x+4=6
-4x + 1 = 6
-4x = 5
dividing through the equation by -4
-4x/-4 = 5/-4
x = -5/4